How much 50% antifreeze solution and 40% antifreeze solution should be combined to give 50 gallons of 46% antifreeze solution?
Found 2 solutions by stanbon, ikleyn:
Answer by stanbon(75887)

How much 50% antifreeze solution and 40% antifreeze solution should be combined to give 50 gallons of 46% antifreeze solution?
-----------------
Equation:
active + active = active
-----
0.50x + 0.40(50-x) = 0.46*50
-----------------
50x + 40*50 - 40x = 46*50
-----------------
10x = 6*50
x = 30 gallons (amt. of 50% solution needed)
50-x = 20 gallons (amt. of 40% solution needed)
-------------
Cheers,
Stan H.
---------------

Answer by ikleyn(52784)

.
For the solution of a similar problem see the lesson  Solving typical word problems on mixtures for solutions  in this site  (Problem 1)>.

Read it attentively and then substitute yours data.